# From Imitation to Refinement Residual RL for Precise Visual Assembly

### v1

@@ 4.1 Improving Imitation Learning with Online Residual Learning
Table 1: Top BC-trained MLPs without chunking (MLP-S) cannot perform any of the tasks, and Diffusion Policies (DP) generally outperform MLPs with chunking (MLP-C). Bottom Training our proposed residual policies with RL on top of frozen diffusion policies performs the best among all evaluated fine-tuning techniques.
Figure 4: Comparison of distilled performance from BC and RL-based teacher.
Fig. 3 shows qualitative examples of common failure modes made by BC policies that are regularly corrected by the RL policy. For instance, a common BC policy error is to push down before the leg is aligned with the hole. This often results in a shift of the object in the grasp, which causes the policy to diverge due to the out-of-distribution grasp pose. We observe the residual policy regularly corrects these errors by performing small sideways translations while canceling out premature downward motions, typically only allowing the leg to be inserted once its properly above the hole. We also find the residual policy is better at performing initial grasps that allow accurate downstream alignment between the grasped object and the receptacle. The BC policies, on the other hand, more often grasp the object at angles that make insertion more difficult.
In addition to improved success rates, we also observed quite different training dynamics across each method. First, direct MLP-C fine-tuning proved unstable and required KL-regularization to avoid collapse. Second, the trained DP produced actions with low variance, even with $\eta=1$ in the DDIM sampler [36], inhibiting Q-learning and constraining the potential for policy improvement. Finally, residual policy training was quite stable, likely because it is constrained to operate on a local scale, which prevents large deviations that can make RL unstable [37, 38].
4.2 Distillation Performance from Synthetic RL Data
Next, we study how synthetic RL data quantity and quality impact the performance of distilled vision-based policies. First, we find that distilling trajectories from the RL agent performs better than training directly on the 50 demonstrations. The RL-distilled vision-based policy reached 73%, outperforming the 50% achieved by training the vision policy directly on human demos. However, we also observe a performance gap between the RL-trained teacher (95%) and distilled student policy (73%). We consider whether this gap may be caused by training the student to operate on images. Upon examination, we find that distilling the same number of teacher rollouts into an image-based student and a state-based student results in comparable performance.
Figure 5: BC distillation scaling with dataset size.
This leads us to conclude that the change in modality is not the primary source of the performance gap. Therefore, we also examine the impact of the distillation dataset size. Here, we scale up the number of state-based rollouts from the trained RL policy and distill these to a state-based student. In Fig. 5, we see that performance increases with more data but still does not reach the performance achieved by the teacher policy, with a gap of 18 percentage points between the best student policy (77% success rate) and the teacher policy (95% success rate) at 10k trajectories. Nevertheless, the improved performance obtained by training with more data highlights the advantage of using simulation for obtaining large-scale synthetic datasets.
4.3 Real-World Performance
Finally, we evaluate the real-world performance of a sim-to-real policy trained on a mixture of a few (10/40) real-world demonstrations and simulation data generated by the trained residual RL policy. We compare the co-trained policy to a baseline model trained solely on real-world demonstrations. We compare the success rates achieved by each policy on two sets of 10 trials for the one_leg task. In the first set, we randomize part poses, while in the second set, we randomize obstacle poses.
The results, presented in Tab. 2, show that incorporating simulation data improves real-world performance (e.g., increasing task completion rate from 20-30% to 50-60%). Qualitatively, the sim-to-real policy exhibits smoother behavior and makes fewer erratic movements that might exceed the robot’s physical limits. Fig. 6 shows examples of successful and unsuccessful task attempts, and more examples of real-world task execution can be found on the website.
See Sec. E.2 for further qualitative experiments probing sim2real transfer and generalizations available from synthetic data.
@@ 40 Real + 350 Sim 10/10 10/10 9/10 10/10 6/10 7/10 5/10 6/10 5/10 6/10
Table 2: We compare the impact of combining real-world demonstrations with simulation trajectories obtained by rolling our RL-trained residual policies. We find that co-training with both real and synthetic data leads to improved motion quality and success rate on the one_leg task.
@@
Figure 6: (A) Examples of successful real world assembly from RGB. Co-training with simulation data reduces jerkiness and improves insertion robustness, due to data from simulation containing higher diversity of initial part poses and insertion locations (see Table 2). (B) Representative failures, primarily due to overfitting to a particular corner location and difficulty adjusting the insertion angle/position when grasps lead to unfamiliar in-hand part poses.
5 Related Works
@@
The approach in [16, 22] studied how to cast diffusion de-noising as a Markov Decision Process, enabling preference-aligned image generation with policy gradient RL. However, this method is unstable and hard to tune. Other ways to combine diffusion architectures with RL include Q-function-based importance sampling [24], advantage weighted regression [25], or changing the objective into a supervised learning problem with return conditioning [26, 27, 39]. Some have also explored augmenting the de-noising training objective with a Q-function maximization objective [28] and iteratively changing the dataset with Q-functions [40]. Recent work developed techniques for training diffusion policies from scratch [23], leveraging combinations of unsupervised clustering and Q-learning to encourage multi-modal behavior discovery. Our method avoids such complexity involved with directly optimizing diffusion models by instead using standard PPO to train simple residual policies that locally correct for errors made by the base policy.
@@
Learning corrective residual components in conjunction with learned or non-learned “base” models has seen widespread success in robotics. Common frameworks include learning residual policies that correct for errors made by a nominal behavior policy [17, 18, 41, 19, 42, 43, 44, 45] and combining learned components to correct for inaccuracies in analytical models for physical dynamics [46, 47, 48] or sensor observations [49]. Residual policies have been used in insertion applications [50] and recent work has applied residual policy learning to the same FurnitureBench task suite we study in this paper [8]. Their approach uses the residual component to model online human-provided corrections via supervised learning, whereas we train our residual policy from scratch with RL using task rewards in simulation.
6 Limitations and Conclusion

### v3

@@ II Problem Setup and Approach Overview
TABLE I: Top BC-trained MLPs without chunking (MLP-S) cannot perform any of the tasks, and Diffusion Policies (DP) generally outperform MLPs with chunking (MLP-C). Bottom Training our proposed residual policies with RL on top of frozen DP performs the best among the evaluated fine-tuning techniques. The best checkpoint is chosen for all methods, and we average over 1024 rollouts to get the success rate.
III Methods
III-A Imitation Learning
For each task, we collect a dataset of 50 demonstrations in simulation by teleoperating the robot, $\mathcal{D}_{\text{sim}}:\ \{\tau_{1},...,\tau_{N}\}$ to train a “base” policy that serves as the starting point for RL training. Each trajectory, $\tau$, contains system states $s_{t}$, and robot actions $a_{t}$, i.e., $\tau_{i}=\{(s_{t},a_{1}),...,(s_{T},a_{T})\}$, with $T$ being the trajectory length.
$\mathcal{D}_{\text{sim}}$ is used to train base policy $\pi_{\text{base}}$ with Behavior Cloning (BC), i.e., $\max_{\pi}\ \mathbb{E}_{(a_{t},s_{t})\sim\mathcal{D}_{\text{sim}}}\left[\log\pi(a_{t}|s_{t})\right]$. We use a Diffusion Policy [18] as the base model, which has shown strong empirical performance in handling difficult manipulation tasks with relatively small datasets [21]. Consistent with recent advancements [18, 6], our policy framework enhances its performance by predicting multiple future actions in chunks instead of individual actions at each timestep. We denote the length of future action sequences predicted by the policy as $T_{a}$, the output as $\mathbf{a_{t}}=[a_{t},...,a_{t+T_{a}}]$. When predicting an action chunk $\mathbf{a_{t}}$ of length $T_{a}$, we only execute a subset $[a_{t},...,a_{t+T_{\text{exec}}}]$, with execution horizon $T_{\text{exec}}\leq T_{a}$. We typically use $T_{a}=32$ and $T_{\text{exec}}=8$.
III-B Online Reinforcement Learning with Residual Policies
Figure 3: Sim-to-real pipeline. (1) Beginning with a policy trained with BC in simulation, (2) we train residual policies with RL and sparse rewards. (3) We then distill the resulting behaviors into a policy operating from RGB images. (4) By combining synthetic data with a small set of real demonstrations, (5) we deploy RGB-based policies in the real world.
Given the initial base policy $\pi_{\text{base}}$, we aim to improve it using RL. However, directly fine-tuning diffusion models with RL is an active area of research [55, 56, 46, 57, 47], made difficult because of the multi-step inference process and unavailability of the policy action log-probabilities. Another category of methods upweight high-quality model outputs via importance sampling, return conditioning, or augmenting the original de-noising objective with a loss term for maximizing a Q-function [58, 59, 60, 17, 61]. However, these methods mainly enable better extraction of high-quality behavior in the data, while we are more concerned with learning new, corrective behaviors. Action chunks can also make optimization with policy gradients more difficult. This is partly due to chunking increasing the effective action space (e.g., chunks of 8 actions increase the dimension 8$\times$), which we find to increase RL training instability (see Sec. IV-A). This means fine-tuning other popular BC architectures, like the Action-Chunked Transformer (ACT) [6], also brings technical challenges [15]. Even with successful RL fine-tuning, the improved policy will still be a trajectory planner lacking in closed-loop reactive control (Fig. 4). Finally, recent work has shown that fine-tuning large pre-trained models can lead to forgetting of capabilities if the agent does not visit states seen during pre-training frequently enough [62].
We side-step these complications by training a residual Gaussian Multi-Layer Perceptron (MLP) [63] policy $\pi_{\text{res}}$. This policy takes as input both the system state and the action predicted by the diffusion policy $\pi_{\text{base}}$ and produces an “action correction” that modifies the action as $a_{t}=a_{t}^{\text{base}}+\alpha\cdot a_{t}^{\text{res}}$, with $\alpha\leq 1$ being a coefficient controlling what scale the residual policy operates on. We find that the policy is not sensitive to $\alpha$ in isolation, but rather the resulting exploration noise, $\alpha\mathcal{N}(0,\sigma^{\text{init std}}I)$, see App. I-C for more analysis. We denote the resulting combined policy $\boldsymbol{\pi}$. This decoupling has several advantages. First, we can regard the diffusion model as a black box and not change its parameters [48]. This decoupling also means that ACT (see App. I-B) and other BC methods can serve as the base model without changing the method – an advantage of residual policy learning emphasized by prior robotics work [48, 64]. This also allows different prediction horizons $T_{a}$ for the two policies. This flexibility in action horizon is helpful as most RL algorithms optimize single-action policies and allow the residual to be fully closed-loop. Finally, it removes the need to explicitly regularize the fine-tuned policy to stay near the pre-trained policy, which is often necessary to achieve stable optimization [65, 66].
Given the above, we train the $\pi_{\text{res}}$ using standard PPO [26] with the alternation that we augment the state space with the base action. The base model observes the current state $s_{t}$ and outputs a chunk of actions $\mathbf{a_{t}}{}$. For each action in the chunk, we concatenate it with the current observation $s_{t+i}^{\text{res}}=[s_{t+i},a^{\text{base}}_{t+i}]$, and predict the correction $a_{t+i}^{\text{res}}\sim\pi^{\text{res}}(\cdot|s_{t+i}^{\text{res}})$ for $i=0,...,T_{a}-1$. Besides the technical convenience, one can look at residual policy as a way to incorporate an inductive bias that the policy should mostly learn locally corrective actions [64]. This is motivated by our qualitative observations that most of the base BC policy failures are due to slight imprecision near “bottleneck” transitions when performing skills like grasping and insertion (e.g., LABEL:fig:overview (Left)).
III-C Sim-to-real pipeline
Ultimately, we are interested in what capabilities our method can enable on real robots. To that end, we devise a simple sim-to-real pipeline to transfer state-based policies trained in simulation to RGB-based policies deployable on physical hardware. We present the general pipeline in Fig. 3. Once the RL training has converged, the combined policy has attained closed-loop correction behavior. We aim to distill this enhanced performance from the state-based policy into a vision-based policy that operates from RGB images. Following the established teacher-student distillation paradigm [31, 67, 68], we generate a dataset of successful trajectories, $\mathcal{D}_{\text{synth}}$, where the environment’s observations, $o_{t}$, replace the states, i.e., $\mathcal{D}_{\text{synth}}=\{\tau_{\text{synth},1},...,\tau_{\text{synth},N}\}$ and $\tau_{\text{synth},i}=\{(o_{1},a_{1}),...,(o_{T},a_{T})\}$. For real-world transfer, we enhance the synthetic dataset $\mathcal{D}_{\text{synth}}$ by re-rendering its trajectories in Isaac Sim [69]. This process improves image quality and introduces variability in environmental conditions such as object and table colors, textures, lighting, and camera perspectives, many of which cannot easily be done with standard image augmentation techniques. We denote this refined dataset as $\mathcal{D}_{\text{synth-render}}$. To ease the difficulty of zero-shot sim-to-real with RGB images, we opt for a co-training approach, integrating the synthetic dataset with a small set of real-world task demonstrations, $\mathcal{D}_{\text{real}}$, which similarly comprises only environmental observations without ground truth poses. This combined dataset, $\mathcal{D}_{\text{real}}\cup\mathcal{D}_{\text{synth-render}}$, is used to train the final student policy with BC.
IV Experiments and Results
@@ 40 Real + 350 Sim 10/10 10/10 9/10 10/10 6/10 7/10 5/10 6/10 5/10 6/10
TABLE II: We investigate the effect of combining real-world demonstrations with simulation trajectories from our RL-trained residual policies. Co-training with real and synthetic data improves motion quality and success rate on the one_leg task.
Figure 4: Perturbations highlight ResiP’s resilience: 12% drop vs. 19-26% for chunked methods. See App. I-A for more detailed discussion.
First, we investigate the requirement of complex BC methods and the impact of our residual reinforcement learning approach on improving the success rates of policies trained with imitation learning in Sec. IV-A. Next, we study the performance of downstream distillation from synthetic RL data, examining the relationship between the quantity and quality of the data and the resulting performance of the distilled vision-based policies in Sec. IV-B. Finally, we apply our approach to real robot hardware, enabling precise assembly tasks on a physical robot directly from vision input in Sec. IV-C.
IV-A Improving Imitation Learning with Residual Learning
Why Action Chunking and Diffusion Policies?
Simple feed-forward MLPs of modest size have shown impressive performance in many domains when trained with RL [68, 67, 30], and offer a natural starting point for RL fine-tuning after BC pre-training. However, the standard MLP policies trained to directly output single action control instead of a trajectory plan through an action chunk (MLP-S) fail across all tasks we consider. Therefore, we also trained MLP policies with action chunking (MLP-C). When we introduce chunking, MLP performance improves drastically, as shown in Tab. I. However, we also find that the more complex Diffusion Policy (DP) architecture generally outperforms MLPs, especially in tasks of intermediate difficulty. For example, an improvement from 10% success rate to 26% for the one_leg task on medium randomness makes subsequent fine-tuning far easier.
In one case, lamp on low randomness, MLP-C outperformed DP. In qualitative evaluations, we find that DP has smoother and faster actions, which is generally beneficial. Still, it seems to hurt performance in this case, as it tends to retract before the gripper fully grasps the lamp base. We also find that all methods struggle with the most challenging tasks, on which MLP-C and DP both achieve less than 5% success rate, indicating that there is still room for improvement in BC methods. The peg-in-hole task, despite its relatively short horizon of $\sim$100 timesteps, proved particularly challenging for BC methods. This task involves a $\sim$0.2 mm tolerance insertion, resulting in a 5% success rate. This poor performance on a short yet precise task lends credence to the hypothesis that BC methods are ill-equipped to handle high-precision requirements.
Another interesting finding is that the DP is significantly more robust to noise, an important property for an adequate base policy for residual fine-tuning. In contrast, we observed performance collapse when using MLP-C as the base, further reinforcing the utility of using powerful generative models for BC pre-training. See more analysis in App. I-B.
Online Policy Fine-Tuning Comparison
Next, we investigate the effectiveness of our residual RL approach, ResiP, in improving the performance of policies trained with BC. We train residual policies using PPO [26] on top of the diffusion policy and compare it to two baselines – (1) directly fine-tune the MLP-C with PPO while regarding every action chunk as one concatenated action and (2) a chunked version of the IDQL algorithm [58]. In the IDQL algorithm, multiple action chunks are sampled from the diffusion policy, and one chooses the action chunk to execute based on its Q-value. We have found that learning a good Q-value estimator from offline data alone, as proposed in IDQL, is challenging with only 50 demonstrations. Therefore, we learn it in on-policy matter similar to [70].
The results in Tab. I show that our residual RL approach significantly improves success rates over BC and outperforms alternative RL fine-tuning methods. For tasks with lower initial randomization, such as the one_leg task, the residual policy increases the success rate from 54% to 98%. Even more drastically, for the peg-in-hole task, ResiP improved the success rate from 5% to 99%. However, we observe performance saturates at lower success rates (e.g., $\sim$70%) for tasks with higher part randomization. We hypothesize that the base model’s performance limits the residual policy as it is designed to act locally, constrained by the base.
Fig. 5 qualitatively shows the most common failure mode of the BC policies. In LABEL:fig:overview (Left), a specific failure mode and how the RL policy overcomes it is further exemplified. For instance, a typical BC policy error is to push down before the leg is aligned with the hole. This often results in a shift of the object in the grasp, which causes the policy to diverge due to the out-of-distribution grasp pose. The residual policy reliably corrects these errors by performing small sideways translations while canceling premature downward motions. It typically only allows the leg to be inserted once it is appropriately above the hole. We also find the residual policy is better at performing initial grasps that allow accurate downstream alignment between the grasped object and the receptacle. On the other hand, the BC policies more often grasp the object at angles that make insertion more difficult. In Fig. 6, we qualitatively analyze this and find that the residual makes the precise insertion part of the task significantly more reliable.
Figure 5: Examples of common failures in tasks that require precise part alignments. All of these failures typically only require small corrections to ameliorate, making the residual framework well-suited to improve on the nominal policy.
Surprisingly, the residual RL procedure also leads to the emergence of qualitatively different behaviors. We refer to the website’s uncut videos of rollouts for the pre-trained and fine-tuned policies. In these videos, we observe that the fine-tuned policy has learned new grasps of the table foot that were not demonstrated, as well as more persistent retrying behavior and recovery.
In addition to improved success rates, we observed different training dynamics across each method. First, direct MLP-C fine-tuning proved unstable and required KL-regularization to avoid collapse. Second, the trained DP produced actions with low variance, even with $\eta=1$ in the DDIM sampler [71], inhibiting Q-learning and constraining the potential for policy improvement. Finally, residual policy training was very stable, likely because it is constrained to operate on a local scale, which prevents large deviations that can make RL unstable and has a small parameter count compared to the base policy [72, 73].
Figure 6: A qualitative analysis of the failure modes between the pre-trained BC policy and the fine-tuned residual policy for the round_table task on medium randomness. For the pre-trained policy, a large portion of the failures (40%) relates to the table leg’s insertion precision. In the fine-tuned policy, this share dropped to 0%.
What drives residual RL performance?
To further understand what drives the residual RL fine-tuning approach’s performance, we try to decompose it into an on-policy component and a closed-loop, reactive component.
The impact of on-policy learning is probed by comparing learning the task by scaling the purely offline data versus learning online with DAgger [74]. The performance implications of distillation are further addressed in Sec. IV-B. We find that when applying DAgger for the same action-chunked diffusion policy, we can improve performance significantly over the purely offline case, as shown in LABEL:fig:overview (right) and Fig. 7. However, a gap remains between the residual RL and DAgger policy of 9% and 23% for the one_leg and round_table tasks, respectively.
On the other hand, we also find the residual policy’s closed-loop, single-step nature helpful. First, when trying to learn a residual policy for whole chunks, the resulting policy learns significantly more slowly, and the performance saturates at a lower level than that of the full closed-loop policy. See App. I-A for details. Furthermore, as shown in Fig. 4, the effect of fully closed-loop control becomes clear when we introduce randomly sampled force perturbations to the episode rollouts. Performance for the ResiP policy drops 12 percentage points, compared to 19-26 percentage points for alternative chunked methods (BC, DAgger, chunked residual).
Figure 7: When scaling up the offline dataset for the harder round_table task, the same trends emerge as for one_leg (seen in LABEL:fig:overview (right)), but with saturation happening at a lower success rate of 56%. Again, correcting for distribution shifts with DAgger improves performance but saturates at 71%.
IV-B Distillation Performance from Synthetic RL Data
Next, we study how synthetic RL data quantity and quality impact the performance of distilled vision-based policies. First, we find that distilling trajectories from the RL agent performs better than training directly on the 50 demonstrations. A vision-based policy distilled from the RL-trained teacher reached 73% on one_leg, outperforming the 50% achieved by training the vision policy directly on human demos, see Fig. 8. However, we also observe a performance gap between the RL-trained teacher (98%) and distilled vision-based student policy (73%). We consider whether this gap may be caused by training the student to operate on images. Upon examination, we find that distilling the same number of teacher rollouts into an image-based student and a state-based student results in comparable performance. This leads us to conclude that the change in modality is not the primary source of the performance gap.
Figure 8: Comparison of distilled performance from BC and RL-based teacher. On the left, BC policies trained on the same set of demonstrations achieve comparable performance for state and image observations. On the right, the vision-based student policy has a performance gap similar to the state-based teacher for state-to-state distillation.
Therefore, we also examine the impact of the distillation dataset size. Here, we scale up the number of state-based rollouts from the trained RL policy and distill these to a state-based student. In LABEL:fig:overview (Right), we see that performance increases with more data but still does not reach the performance achieved by the teacher policy, with a gap of 20 percentage points between the best student policy (80% success rate) and the teacher policy (98% success rate) at 100k trajectories (with only a minor improvement from 78% at 10k trajectories). The same trend is evident in Fig. 7 (though the saturation occurs earlier). Nevertheless, the improved performance obtained by training with more data highlights the advantage of using simulation for obtaining large-scale synthetic datasets.
DAgger [1] demonstrates significantly greater sample and computational efficiency than pure BC. As shown in LABEL:fig:overview (right), DAgger achieves superior performance to BC trained on 50 human demos in just $\sim$10k gradient steps (with 800 rollouts) and surpasses our best BC distillation results at $\sim$20k steps (with $\sim$1.5k demonstrations). These results underscore the advantages of interactive learning over passive offline learning and point towards effective online distillation techniques to help close the gap between the state-based RL-trained experts and the distilled real-world policies deployed on real hardware (see Sec. IV-C).
IV-C Real-World Application
Figure 9: Examples of photo-realistic rendering and domain randomization for sim-to-real transfer. In the first row, we show our nominal experiment setting of rendering parts in the original white color while varying the position, brightness, and hue of the lighting in the scene. In the bottom row, we also introduce variations in the part colors to see how that can help the final policy adapt to unseen part appearances without collecting any additional data.
@@ IV-C Real-World Application
Figure 10: (A) Examples of successful real world assembly from RGB. Co-training with simulation data reduces jerkiness and improves insertion robustness by containing a higher diversity of part poses and insertion locations (see Table II). (B) Example failure: difficulty adjusting the insertion angle/position when grasps lead to unseen in-hand part poses.
Finally, we evaluate the real-world performance of a sim-to-real policy trained on a mixture of a few (10/40) real-world demonstrations and simulation data generated by the trained residual RL policy. We compare the co-trained policy to a baseline model trained only on real-world demonstrations. We compare the success rates achieved by each policy on two sets of 10 trials for the one_leg task. In the first set, we randomize part poses, while in the second set, we randomize obstacle poses (i.e., insertion location in the workspace).
The results, presented in Tab. II, show that incorporating simulation data improves real-world performance (e.g., increasing task completion rate from 20-30% to 50-60%). Qualitatively, the sim-to-real policy exhibits smoother behavior and makes fewer erratic movements that might exceed the robot’s physical limits. Fig. 10 shows examples of successful and unsuccessful task attempts.
To further probe the sim-to-real transfer and potential for acquiring new capabilities with the help of the simulator, we create a task variation where the part colors are changed from black to white. When rolling out the policy trained on real demos of white parts, the robot exhibited erratic behavior that caused the hardware to reach velocity limits on every trial we ran, as shown in Fig. 11 (A). When including synthetic data rendered with parts in black, the resulting policy can perform the task again (see Fig. 11). The resulting performance was still inferior to the performance on white parts, which motivates further work on closing the sim-to-real gap.
Figure 11: (A) When changing the part appearances from white to black, the policy trained on real data only seizes to function. The behavior becomes erratic, and all trials ended with the robot hitting a velocity limit. (B) When mixing in synthetic data with more diverse colors (see Fig. 9), the policy regains the ability to complete the task, though with lower performance than for white parts.
V Related Works
@@
The approach in [46, 56] studied how to cast diffusion de-noising as a Markov Decision Process, enabling preference-aligned image generation with policy gradient RL. However, this method is unstable and difficult to tune. [47] introduced a framework for direct diffusion policy fine-tuning, which avoids this instability, but the method is architecture-specific and does not introduce closed-loop control. Other ways to combine diffusion architectures with RL include Q-function-based importance sampling [58], advantage weighted regression [59], or changing the objective into a supervised learning problem with return conditioning [60, 17, 16]. Some have also explored augmenting the de-noising training objective with a Q-function maximization objective [61] and iteratively changing the dataset with Q-functions [75]. Recent work developed techniques for training diffusion policies from scratch [57], leveraging unsupervised clustering and Q-learning ensembles to encourage multi-modal behavior discovery. Our method avoids such complexity involved with directly optimizing diffusion models by using standard PPO to train simple residual policies that correct for errors made by the base policy.
@@
Learning corrective residual components in conjunction with learned or non-learned “base” models has been widely successful in robotics. Common frameworks include learning residual policies that correct for errors made by a nominal behavior policy [48, 64, 76, 77, 78, 79, 80, 81] and combining learned components to correct for inaccuracies in analytical models for physical dynamics [82, 83, 84] or sensor observations [85]. Residual policies have been used in insertion applications [86], and recent work has applied residual policy learning to the same FurnitureBench task suite we study in this paper [87]. Their approach uses the residual component to model online human-provided corrections via supervised learning, whereas we train our residual policy from scratch with RL using task rewards in simulation.
VI Limitations and Conclusion